# In-Model Merging for Enhancing the Robustness of Medical Imaging Classification Models

### v1

@@ 3.3 Analyses
Ablation Study, and Merge with different merge weights $\alpha$ / probabilities $p$. From Fig. 2(c), we can observe from the perspective of either different merge weights $\alpha$ or probabilities $p$, the line plots show a stable trend and highly above the baseline model (trained with the same number of epochs of InMerge), also serving as the ablation study of w and w/o InMerge.
In-model Merging on different CNN architectures. We also examine the generality of In-model Merging on different models. For VGG16 model ($L_{s}=$10), the average AUC of In-model Merging surpasses the baseline model (0.817 VS. 0.807). Similar results are shown on ResNet34 ($L_{s}=$30), the model performance gained from 0.792 to 0.796.
(a) Layer thresholds.
(b) Different similarities.
(c) Merge weight + prob.
@@
Figure 2: The sensitivity of different merged layers (a), merge with different similarities (b), merge with different merge weights and merge probabilities (c).
@@ 4 Conclusion
We propose In-model Merging, where the merging process occurs within one model and enhances the model robustness with minimal training computations and no additional computations in inference. By selectively merging similar kernels in deeper layers, the method enhances feature representation while preserving low-level features for robustness. Our experiments confirm the effectiveness of in-model merging in improving model performance and robustness. Results across multiple medical datasets show improved generalization. This suggests that in-model merging is a practical enhancement for neural networks with potential for further exploration. Future research may focus on exploring InMerge in different network architectures, such as Transformers [17] or Mamba [9].

### v2

@@ 4.3.1 Ablation: Merging with Different Weights $\alpha$ and Probabilities $p$
From Fig. 2, we can observe from the perspective of either different merge weights $\alpha$ or probabilities $p$, the line plots show a stable trend and highly above the baseline model (VGG19 trained with the same number of epochs, but without InMerge operations), also serving as the ablation study of w and w/o InMerge.
4.3.2 Merging Layer $L_{s}$ Sensitivity
@@
Figure 3: (a) The sensitivity of different merged layers; (b) Merging with different similarities.
Fig. 3 (a) shows performance variations across merging different layers ($\geq$ Layer 10 means any layer below 10th layer is not considered for merging). A clear trend shows where only including deeper layers (Layer 10 and Layer 15) outperform including from shallower ones (Layer 0 and Layer 5) in most categories. This highlights the importance of deep feature representations for disease identification. Notably, conditions such as Emphysema, Cardio., and Pneumo. show substantial improvements when merging deeper layers, suggesting the possibility that these conditions rely more on low-level abstract features, so they are sensitive to layer change. In contrast, some categories, such as Infiltration and Consolidation, exhibit relatively stable or even slightly better performance at shallower layers, possibly due to their reliance on low-level texture features. For the setup of layer 0 and 5, the model performances are very different from layer 10, but models with $\geq$ layer 10 and 15 performance are similar, showing the general stability when layers go deeper. These findings empirically prove the idea that deeper feature merging is crucial for model merging strategies, while also indicating that the optimal layer depth for different diseases may vary, which is valuable for task-specific model merging setup.
4.3.3 Merging with Different Similarities $\tau$
Fig. 3 (b) compares model performance under varying similarity thresholds for in-model kernel merging. A clear trend emerges where higher similarity thresholds (e.g., $\geq$ Sim 0.3) generally lead to stronger performance than the thresholds 0.1 or Dissimilar condition (we show dissimilar threshold 0.5 in the figure as well), where more dissimilar kernels are merged, indicating that merging dissimilar kernels disrupts critical feature representations. Interestingly, for Effusion and Edema, the performance difference between different similarity settings is relatively small, suggesting that these conditions may be less sensitive to in-model merging. These results highlight the importance of selective merging strategies, where preserving high-similarity features is beneficial while merging dissimilar kernels can degrade performance.
4.3.4 In-model Merging on Different CNN Architectures
We also examine the generality of In-model Merging on different models. For VGG16 model ($L_{s}=$10), the average AUC of In-model Merging surpasses the baseline model (0.817 VS. 0.807). Similar results are shown on ResNet34 ($L_{s}=$30), the model performance gained from 0.792 to 0.796.
@@ 5 Conclusion
We propose In-model Merging, where the merging process occurs within one model and enhances the model robustness with minimal training computations and no additional computations in inference. By selectively merging similar kernels in deeper layers, the method enhances feature representation while preserving low-level features for robustness. Our experiments confirm the effectiveness of in-model merging in improving model performance and robustness. Results across multiple medical datasets show improved generalization. This suggests that in-model merging is a practical enhancement for neural networks with potential for further exploration. Future research may focus on exploring InMerge in different network architectures, such as Transformers [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] or Mamba [Gu and Dao(2023)].